# Predictive Linear Online Tracking for Unknown Targets

### v1

@@ B.3 Perturbation Analysis
Appendix C Proofs for Prediction
@@
C.1 Regret guarantees for one-step ahead prediction
@@
0:  forgetting factor $\gamma\in(0,1)$, initial regularization $\epsilon$
  Initialize: $P_{-1}=\epsilon I_{np}$, $\hat{S}_{0}\in\mathcal{S}$;
@@
To bound the regret of RLS, we adapt the proof of Yuan and Lamperski [2020] while keeping track of all quantities of interest, e.g. the system dimension, logarithmic terms etc., and while working with matrices instead of vectors.
Theorem C.1.
@@
$\beta_{1}=2\sqrt{n}M\left(\frac{\epsilon+pD^{2}_{r}}{1-\gamma}\right),\,\beta_{2}=np(1+\sqrt{p}M)^{2}D^{2}_{r},\,$ (23)
$\beta_{3}=\beta_{2},\,\beta_{4}=2\gamma\epsilon nM^{2}+\beta_{2}\log\frac{\epsilon+pD^{2}_{r}}{\epsilon}.$
@@
where we used the identity $a^{\top}b=\mathrm{tr}(a^{\top}b)=\mathrm{tr}(ba^{\top})$. Invoking Lemma C.2 and Lemma C.3, we have
$\mathcal{R}_{\mathrm{pred}}(S_{0:T})$ $\leq 2\sqrt{p}M(\epsilon+\frac{pD^{2}_{r}}{1-\gamma})V_{T}+\frac{1}{2}\max_{t\leq T}\|r_{t}-\hat{S}_{t}z_{t-1}\|^{2}\sum_{t=0}^{T}z_{t-1}^{\top}P^{-1}_{t}z_{t-1}+2\gamma\epsilon nM^{2}.$
The result follows from Lemma C.5 and the fact that $\|r_{t}\|\leq D_{r}$, $\|\hat{S}_{t}z_{t-1}\|\leq M\|z_{t-1}\|\leq M\sqrt{p}D_{r}$ since $z_{t}$ a concatenation of $p$ vectors. As a result, we can bound the error $\|r_{t}-\hat{S}_{t}z_{t-1}\|^{2}$ by $(1+\sqrt{p}M)^{2}D^{2}_{r}$ for all $t\leq T$. To simplify the final bound we use $T+1\leq 2T$. ∎
The following technical lemmas are auxiliary results towards proving Theorem C.1. Lemmas C.2C.3 control the growth of the first order (gradient) term in the regret. Lemma C.4 upper bounds matrix $P_{t}$ while Lemma C.5 contains a standard elliptical potential bound, tailored to the forgetting factor case.
Lemma C.2 (Gradient Inner Product).
Consider the conditions of Theorem C.1. Let $\hat{S}_{t}$ be the RLS estimate at time $t$ and $S_{t},S_{t+1}$ be any arbitrary matrices that satisfy the constraints, i.e., $S_{t},\,S_{t+1}\in\mathcal{S}$. We have
@@
$\leq\|\hat{S}_{t}-S_{t}\|^{2}_{F,P_{t}}-\|\hat{S}_{t+1}-S_{t+1}\|^{2}_{F,P_{t}}+\|r_{t}-\hat{S}_{t}z_{t-1}\|^{2}z^{\top}_{t-1}P^{-1}_{t}z_{t-1}+4\sqrt{p}M(\epsilon+\frac{pD^{2}_{r}}{1-\gamma})\|S_{t+1}-S_{t}\|_{F}$
@@
$\geq\|\hat{S}_{t+1}-S_{t+1}\|^{2}_{F,P_{t}}-4\sqrt{n}M(\epsilon+\frac{pD^{2}_{r}}{1-\gamma})\|S_{t+1}-S_{t}\|_{F},$
where we used $\|\hat{S}_{t+1}-S_{t+1}\|_{F}\leq\sqrt{\mathrm{rank}(\hat{S}_{t+1}-S_{t+1})}\|\hat{S}_{t+1}-S_{t+1}\|$, $\mathrm{rank}(\hat{S}_{t+1}-S_{t+1})\leq n$, $\|\hat{S}_{t+1}-S_{t+1}\|\leq 2M$ and Lemma C.4. Combining the above two inequalities gives us the result. ∎
Lemma C.3 (Telescoping Series).
Consider the conditions of Theorem C.1.
$\sum_{t=0}^{T}\|\hat{S}_{t}-S_{t}\|^{2}_{F,P_{t}}-\|\hat{S}_{t+1}-S_{t+1}\|^{2}_{F,P_{t}}\leq\sum_{t=0}^{T}\mathrm{tr}((\hat{S}_{t}-S_{t})z_{t-1}z_{t-1}^{\top}(\hat{S}_{t}-S_{t})^{\top})+\gamma\epsilon\,\mathrm{tr}((\hat{S}_{0}-S_{0})(\hat{S}_{0}-S_{0})^{\top}).$
@@
$\|\hat{S}_{0}-S_{0}\|^{2}_{F,P_{0}}=\mathrm{tr}((\hat{S}_{0}-S_{0})(z_{-1}z_{-1}^{\top}+\gamma\epsilon I)(\hat{S}_{0}-S_{0})^{\top})$
@@
Lemma C.4 (Design matrix bound).
Consider the conditions of Theorem C.1 with $P_{t}=\gamma P_{t-1}+z_{t-1}z_{t-1}^{\top}$ and $P_{-1}=\epsilon I$. We have
$\|P_{t}\|\leq\epsilon+\frac{pD^{2}_{r}}{1-\gamma}.$
@@
$\|P_{t}\|\leq\gamma\|P_{t-1}\|+pD^{2}_{r}=\frac{1-\gamma^{t}}{1-\gamma}pD^{2}_{r}+\gamma^{t+1}\epsilon\leq\frac{1}{1-\gamma}pD^{2}_{r}+\epsilon$
Lemma C.5 (Forgetting potential lemma).
Consider the conditions of Theorem C.1 with $P_{t}=\gamma P_{t-1}+z_{t-1}z_{t-1}^{\top}$ and $P_{-1}=\epsilon I$. The following upper bound is true
$\sum_{t=0}^{T}z_{t-1}^{\top}P^{-1}_{t}z_{t-1}\leq np\log(\frac{\epsilon+pD^{2}_{r}}{\epsilon(1-\gamma)})+np(T+1)\log\frac{1}{\gamma}.$
@@
$\sum_{t=0}^{T}z_{t}^{\top}P^{-1}_{t}z_{t}\leq\log\det P_{T}-\log\det(\epsilon I_{np})+np(T+1)\log 1/\gamma.$
The final bound follows from the upper bound on $P_{T}$ given in Lemma C.4. ∎
C.2 Regret guarantees for multi-step ahead prediction
Consider now the case of $k-$steps ahead prediction. Let $S_{k-1|-1},\,S_{k|0},\dots,S_{T|T-k}$ be any comparator sequence that satisfies Assumption 2.2. In this case, the regret of $k-$step ahead prediction is defined as
@@ C.2 Regret guarantees for multi-step ahead prediction
Recall that we have $k$ learners updated at non-overlapping time steps. Let $i$, for $i=0,\dots,k-1$, denote the index for the learner. We will decompose the problem into $k$ independent one-step ahead prediction problems and invoke Theorem C.1.
@@
As a consequence of redefining the time axis, the prediction regret of every individual learner $i$ can be bounded using Theorem C.1. Let $V^{k,i}_{T}=\sum_{\tau=0}^{N_{k,i}-1}\|S^{i}_{\tau+1}-S^{i}_{\tau}\|$. Then we obtain
@@
Appendix D Regret of the PLOT algorithm
Let $\Sigma=B^{\top}PB+R$. Consider again the optimal non-causal policy
@@ Appendix D Regret of the PLOT algorithm
By invoking the performance difference lemma, we obtain that the regret is equal to
@@
$=\sum_{t=0}^{T-1}\lVert\hat{q}_{t}(r_{t:t+W|t}-r_{t:t+W})+\delta_{t}(r_{t+W:T})\rVert^{2}_{\Sigma}.$
@@
Lemma D.1 (Truncation term).
@@
The first equality follows by the definition of $\delta_{t}(r_{t+W:T})$ and the fact that $\delta_{t}(r_{t+W:T})=0$, for $t+W>T-1$. By Proposition A.1, we have $\|K_{t}\|\leq c_{0}\rho^{t}$. Hence
@@
Lemma D.2 (Prediction term).
@@
$L_{i}=K_{i}A-K_{i-1}$ $=\Sigma^{-1}B^{\top}(A-BK)^{\top,i}PA-\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}P$
$=\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}((A-BK)^{\top}PA-P)$
$=\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}(A^{\top}PA-K^{\top}B^{\top}PA-P)$
$=\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}(-Q+K^{\top}B^{\top}PA-K^{\top}B^{\top}PA)$
@@
Hence, by Proposition A.1 and the fact that $Q\preceq P$, we also get
@@
Lemma D.3 (Geometric Series).
@@
D.1 Proof of Theorem 4.1
By (25), Lemma D.1, and Lemma D.2, we obtain that
@@
By the perturbation bound in Lemma B.3, we can upper bound $V^{i}_{T}\leq\sqrt{p}M^{2}i^{2}V_{T}$. Hence, by invoking Lemma D.3 we finally obtain
@@
where $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ are given in (23) and $\|\Sigma\|=\|B^{\top}PB+R\|$.
D.2 Proof of Corollary 4.2
@@
Appendix E Simulations
E.1 The Quadrotor Model
To demonstrate the performance of PLOT on a practical use-case we consider the linearized dynamics of the Crazyflie quadrotor Bitcraze [2023b], derived in Beuchat [2019]. The mini quadrotor can be modeled with a continuous-time nonlinear model $g:\mathbb{R}^{9}\times\mathbb{R}^{4}\rightarrow\mathbb{R}^{9}$
@@ E.1 The Quadrotor Model
with ${p}:=[p_{x};p_{y};p_{z}]^{\top}$ defined as the position vector in the inertial frame, ${\psi}:=[\gamma,\beta,\alpha]^{\top}$ as the attitude vector in the same frame and $\gamma$, $\beta$ and $\alpha$ denote the attitude angles, roll, pitch and yaw, respectively. The action ${u}:=[f;{\omega}]^{\top}$ comprises of the total thrust $f$, as well as the angular rate ${\omega}:=[\omega_{x},\omega_{y},\omega_{z}]^{\top}$ in the body frame. We provide a brief derivation of the linearized model for the quadrotor of interest and refer the readers to Beuchat [2019] for a detailed derivation.
@@
where $x_{h}=\boldsymbol{0}\in\mathbb{R}^{9}$ and $u=[mg;0;0;0]^{\top}$ are the hovering position, steady state, and input, and $\delta x:=x-x_{h},\;\delta u:=u-u_{h}$. Calculating the respective Jacobians, and discretizing the resulting dynamics with a sampling time $T_{s}$, the following dynamics are derived
@@
In this section, we first analyze PLOT, showing the effect of the prediction horizon, $W$ on the tracking performance, and how this is reflected in the dynamic regret bound. We also show how the regret bound can be used to adjust the forgetting factor, $\gamma$, to achieve better tracking given the order of the reference target path length. Next, we compare the proposed method with state-of-the-art online control algorithms for the non-stochastic adversarial control setting with both static and dynamic regret bounds. We show how for certain dynamic references the proposed indirect method of PLOT provides better tracking and lower dynamic regret.
E.2 PLOT: Regret for hyperparameter tuning

### v3

@@ C.3 Perturbation Analysis
Appendix D Proofs for Prediction
@@
D.1 Regret guarantees for one-step ahead prediction
@@
0:  forgetting factor $\gamma\in(0,1)$, initial regularization $\varepsilon$
  Initialize: $P_{-1}=\varepsilon I_{np}$, $\hat{S}_{0}\in\mathcal{S}$;
@@
To bound the regret of RLS, we adapt the proof of (Yuan & Lamperski, 2020) while keeping track of all quantities of interest, e.g. the system dimension, logarithmic terms etc., and while working with matrices instead of vectors.
Theorem D.1.
@@
$\beta_{1}=2\sqrt{n}M\left(\frac{\varepsilon+pD^{2}_{r}}{1-\gamma}\right),\,\beta_{2}=np(1+\sqrt{p}M)^{2}D^{2}_{r},$ (23)
$\beta_{3}=\beta_{2},\,\beta_{4}=2\gamma\varepsilon nM^{2}+\beta_{2}\log\frac{\varepsilon+pD^{2}_{r}}{\varepsilon}.$
@@
where we used the identity $a^{\top}b=\mathrm{tr}(a^{\top}b)=\mathrm{tr}(ba^{\top})$. Invoking Lemma D.2 and Lemma D.3, we have
$\mathcal{R}_{\mathrm{pred}}(S_{0:T})$ $\leq 2\sqrt{p}M(\varepsilon+\frac{pD^{2}_{r}}{1-\gamma})V_{T}+\frac{1}{2}\max_{t\leq T}\|r_{t}-\hat{S}_{t}z_{t-1}\|^{2}\sum_{t=0}^{T}z_{t-1}^{\top}P^{-1}_{t}z_{t-1}+2\gamma\varepsilon nM^{2}.$
The result follows from Lemma D.5 and the fact that $\|r_{t}\|\leq D_{r}$, $\|\hat{S}_{t}z_{t-1}\|\leq M\|z_{t-1}\|\leq M\sqrt{p}D_{r}$ since $z_{t}$ a concatenation of $p$ vectors. As a result, we can bound the error $\|r_{t}-\hat{S}_{t}z_{t-1}\|^{2}$ by $(1+\sqrt{p}M)^{2}D^{2}_{r}$ for all $t\leq T$. To simplify the final bound we use $T+1\leq 2T$. ∎
The following technical lemmas are auxiliary results towards proving Theorem D.1. Lemmas D.2D.3 control the growth of the first order (gradient) term in the regret. Lemma D.4 upper bounds matrix $P_{t}$ while Lemma D.5 contains a standard elliptical potential bound, tailored to the forgetting factor case.
Lemma D.2 (Gradient Inner Product).
Consider the conditions of Theorem D.1. Let $\hat{S}_{t}$ be the RLS estimate at time $t$ and $S_{t},S_{t+1}$ be any arbitrary matrices that satisfy the constraints, i.e., $S_{t},\,S_{t+1}\in\mathcal{S}$. We have
@@
$\leq\|\hat{S}_{t}-S_{t}\|^{2}_{F,P_{t}}-\|\hat{S}_{t+1}-S_{t+1}\|^{2}_{F,P_{t}}+\|r_{t}-\hat{S}_{t}z_{t-1}\|^{2}z^{\top}_{t-1}P^{-1}_{t}z_{t-1}+4\sqrt{p}M(\varepsilon+\frac{pD^{2}_{r}}{1-\gamma})\|S_{t+1}-S_{t}\|_{F}$
@@
$\geq\|\hat{S}_{t+1}-S_{t+1}\|^{2}_{F,P_{t}}-4\sqrt{n}M(\varepsilon+\frac{pD^{2}_{r}}{1-\gamma})\|S_{t+1}-S_{t}\|_{F},$
where we used $\|\hat{S}_{t+1}-S_{t+1}\|_{F}\leq\sqrt{\mathrm{rank}(\hat{S}_{t+1}-S_{t+1})}\|\hat{S}_{t+1}-S_{t+1}\|$, $\mathrm{rank}(\hat{S}_{t+1}-S_{t+1})\leq n$, $\|\hat{S}_{t+1}-S_{t+1}\|\leq 2M$ and Lemma D.4. Combining the above two inequalities gives us the result. ∎
Lemma D.3 (Telescoping Series).
Consider the conditions of Theorem D.1.
$\sum_{t=0}^{T}\|\hat{S}_{t}-S_{t}\|^{2}_{F,P_{t}}-\|\hat{S}_{t+1}-S_{t+1}\|^{2}_{F,P_{t}}\leq\sum_{t=0}^{T}\mathrm{tr}((\hat{S}_{t}-S_{t})z_{t-1}z_{t-1}^{\top}(\hat{S}_{t}-S_{t})^{\top})+\gamma\varepsilon\,\mathrm{tr}((\hat{S}_{0}-S_{0})(\hat{S}_{0}-S_{0})^{\top}).$
@@
$\|\hat{S}_{0}-S_{0}\|^{2}_{F,P_{0}}=\mathrm{tr}((\hat{S}_{0}-S_{0})(z_{-1}z_{-1}^{\top}+\gamma\varepsilon I)(\hat{S}_{0}-S_{0})^{\top})$
@@
Lemma D.4 (Design matrix bound).
Consider the conditions of Theorem D.1 with $P_{t}=\gamma P_{t-1}+z_{t-1}z_{t-1}^{\top}$ and $P_{-1}=\varepsilon I$. We have
$\|P_{t}\|\leq\varepsilon+\frac{pD^{2}_{r}}{1-\gamma}.$
@@
$\|P_{t}\|\leq\gamma\|P_{t-1}\|+pD^{2}_{r}=\frac{1-\gamma^{t}}{1-\gamma}pD^{2}_{r}+\gamma^{t+1}\varepsilon\leq\frac{1}{1-\gamma}pD^{2}_{r}+\varepsilon$
Lemma D.5 (Forgetting potential lemma).
Consider the conditions of Theorem D.1 with $P_{t}=\gamma P_{t-1}+z_{t-1}z_{t-1}^{\top}$ and $P_{-1}=\varepsilon I$. The following upper bound is true
$\sum_{t=0}^{T}z_{t-1}^{\top}P^{-1}_{t}z_{t-1}\leq np\log(\frac{\varepsilon+pD^{2}_{r}}{\varepsilon(1-\gamma)})+np(T+1)\log\frac{1}{\gamma}.$
@@
$\sum_{t=0}^{T}z_{t}^{\top}P^{-1}_{t}z_{t}\leq\log\det P_{T}-\log\det(\varepsilon I_{np})+np(T+1)\log 1/\gamma.$
The final bound follows from the upper bound on $P_{T}$ given in Lemma D.4. ∎
D.2 Regret guarantees for multi-step ahead prediction
Consider now the case of $k-$steps ahead prediction. Recall that for any fixed $k=1,\dots,W$, the $k-$step ahead predictor maintains $k$ learners that are updated at non-overlapping intervals. An example for $k=1,2,3$ can be found in Figure 7.
Figure 7: The updates of the $k=1,2,3$ steps ahead predictors for $t=2,\dots,6$. Projections are omitted for simplicity. Each predictor consists of $k$ independent learners that are updated at non-overlapping time steps. For example, for $k=2$, there are two learners and the update for $\hat{S}_{t+2|t}$ is decoupled from the one of $\hat{S}_{t+1|t-1}$.
Let $S_{k-1|-1},\,S_{k|0},\dots,S_{T|T-k}$ be any comparator sequence that satisfies Assumption 2.2. In this case, the regret of $k-$step ahead prediction is defined as
@@ D.2 Regret guarantees for multi-step ahead prediction
Recall that we have $k$ learners updated at non-overlapping time steps. Let $i$, for $i=0,\dots,k-1$, denote the index for the learner. We will decompose the problem into $k$ independent one-step ahead prediction problems and invoke Theorem D.1.
@@
As a consequence of redefining the time axis, the prediction regret of every individual learner $i$ can be bounded using Theorem D.1. Let $V^{k,i}_{T}=\sum_{\tau=0}^{N_{k,i}-1}\|S^{i}_{\tau+1}-S^{i}_{\tau}\|$. Then we obtain
@@
Appendix E Regret of the PLOT algorithm
Let $\Sigma=B^{\top}XB+R$. Consider again the optimal non-causal policy
@@ Appendix E Regret of the PLOT algorithm
By invoking the performance difference lemma (Lemma 4.4), we obtain that the regret is equal to
@@
$=\sum_{t=0}^{T-1}\lVert\hat{q}_{t}(r_{t:t+W|t}-r_{t:t+W}){{-}}\delta_{t}(r_{t+W:T})\rVert^{2}_{\Sigma}.$
@@
Lemma E.1 (Truncation term).
@@
The first equality follows by the definition of $\delta_{t}(r_{t+W:T})$ and the fact that $\delta_{t}(r_{t+W:T})=0$, for $t+W>T-1$. By Proposition B.1, we have $\|K_{t}\|\leq c_{0}\rho^{t}$. Hence
@@
Lemma E.2 (Prediction term).
@@
$L_{i}=K_{i}A-K_{i-1}$ $=\Sigma^{-1}B^{\top}(A-BK)^{\top,i}XA-\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}X$
$=\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}((A-BK)^{\top}XA-X)$
$=\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}(A^{\top}XA-K^{\top}B^{\top}XA-X)$
$=\Sigma^{-1}B^{\top}(A-BK)^{\top,i-1}(-Q+K^{\top}B^{\top}XA-K^{\top}B^{\top}XA)$
@@
Hence, by Proposition B.1 and the fact that $Q\preceq X$, we also get
@@
Lemma E.3 (Geometric Series).
@@
E.1 Proof of Theorem 4.1
By (25), Lemma E.1, and Lemma E.2, we obtain that
@@
By the perturbation bound in Lemma C.3, we can upper bound $V^{i}_{T}\leq\sqrt{p}M^{2}i^{2}V_{T}$. Hence, by invoking Lemma E.3 we finally obtain
@@
where $\beta_{1},\beta_{2},\beta_{3},\beta_{4}$ are given in (23) and $\|\Sigma\|=\|B^{\top}XB+R\|$.
E.2 Proof of Corollary 4.2
@@
Appendix F Simulations
F.1 The Quadrotor Model
To demonstrate the performance of PLOT on a practical use-case we consider the linearized dynamics of the Crazyflie quadrotor (Bitcraze, 2023b), derived in (Beuchat, 2019). The mini quadrotor can be modeled with a continuous-time nonlinear model $g:\mathbb{R}^{9\times 1}\times\mathbb{R}^{4\times 1}\rightarrow\mathbb{R}^{9\times 1}$
@@ F.1 The Quadrotor Model
with ${p}:=[p_{x};p_{y};p_{z}]^{\top}$ defined as the position vector in the inertial frame, ${\psi}:=[\gamma,\beta,\alpha]^{\top}$ as the attitude vector in the same frame and $\gamma$, $\beta$ and $\alpha$ denote the attitude angles, roll, pitch and yaw, respectively. The action ${u}:=[f;{\omega}]^{\top}$ comprises of the total thrust $f$, as well as the angular rate ${\omega}:=[\omega_{x},\omega_{y},\omega_{z}]^{\top}$ in the body frame. We provide a brief derivation of the linearized model for the quadrotor of interest and refer the readers to (Beuchat, 2019) for a detailed derivation.
@@
where $x_{h}=\boldsymbol{0}\in\mathbb{R}^{9\times 1}$ and $u=[mg;0;0;0]^{\top}$ are the hovering position, steady state, and input, and $\delta x:=x-x_{h},\;\delta u:=u-u_{h}$. Calculating the respective Jacobians, and discretizing the resulting dynamics with a sampling time $T_{s}$, the following dynamics are derived
@@
For the PLOT policy, we take $p=1$ for all experiments, and the projection step $8$ in the RLS Algorithm 1 is executed with a quadratic program solver, solving the following problem
$\hat{S}_{t+k|t}=\arg\min_{S}$ $\|S-Y\|_{F,P_{t|t-k}}$
s.t. $\|\text{vec}(S)\|_{\infty}\leq M,$
where $\text{vec}(S)$ is a vector formed by vertically stacking the columns of $S$. Note that, the $\|\cdot\|_{\infty}$ constraint is a conservative approximation of the original problem, resulting in a safe, and conservative estimate of the original projection $\Pi_{\mathcal{S}}^{P_{t|t-k}}$.
In this section, we first analyze PLOT, showing the effect of the prediction horizon, $W$ on the tracking performance, and how this is reflected in the dynamic regret bound. We also show how the regret bound can be used to adjust the forgetting factor, $\gamma$, to achieve better tracking given the order of the reference target path length. Next, we compare the proposed method with state-of-the-art online control algorithms for the non-stochastic adversarial control setting with both static and dynamic regret bounds. We show how for certain dynamic references the proposed indirect method of PLOT provides better tracking and lower dynamic regret, and how for others it can perform worse than direct methods.
F.2 PLOT: Regret for hyperparameter tuning